If Lucy is four times the age of Amy now,
three years ago Lucy was 13 times the Age of Amy, how old is Lucy?
please show me how to figure this out as I am totally lost.
Thanks

Answer by scott8148(5885)

NOW ___ let x=Lucy, so x/4=Amy

3 YEARS AGO ___ Lucy=x-3, Amy=(x/4)-3

"three years ago Lucy was 13 times the Age of Amy" ___ x-3=13((x/4)-3)

x-3=(13/4)x-39